OpenStudy (brittany):

simplify this expression 1/3√144

OpenStudy (anonymous):

4

OpenStudy (anonymous):

12^2 =144

OpenStudy (angela210793):

sqr144=12--->1/3*12=12/3=4

OpenStudy (brittany):

whats the answer

OpenStudy (anonymous):

4...

OpenStudy (anonymous):

\[\left(\frac{1}{3}\right)\sqrt{144}=4 \]

OpenStudy (anonymous):

(1/3)*12=4

OpenStudy (anonymous):

or 1/(3*12)=1/36